# Context Matters: Learning Generalizable Rewards via Calibrated Features

### v2

@@ Abstract
A key challenge in reward learning from human input is that desired agent behavior often changes based on context. Traditional methods typically treat each new context as a separate task with its own reward function. For example, if a previously ignored stove becomes too hot to be around, the robot must learn a new reward from scratch, even though the underlying preference for prioritizing safety over efficiency remains unchanged. We observe that context influences not the underlying preference itself, but rather the saliency—or importance—of reward features. For instance, stove heat affects the importance of the robot’s proximity, yet the human’s safety preference stays the same. Existing multi-task and meta IRL methods learn context-dependent representations implicitly—without distinguishing between preferences and feature importance—resulting in substantial data requirements. Instead, we propose explicitly modeling context-invariant preferences separately from context-dependent feature saliency, creating modular reward representations that adapt to new contexts. To achieve this, we introduce calibrated features—representations that capture contextual effects on feature saliency—and present specialized paired comparison queries that isolate saliency from preference for efficient learning. Experiments with simulated users show our method significantly improves sample efficiency, requiring 10x fewer preference queries than baselines to achieve equivalent reward accuracy, with up to 15% better performance in low-data regimes (5–10 queries). An in-person user study (N=12) demonstrates that participants can effectively teach their unique personal contextual preferences using our method, enabling more adaptable and personalized reward learning.
1 Introduction
Imagine a robot helping you in the kitchen. When the stove is off, the robot can pass freely above it to reach the counter. When the stove starts heating up, the robot may keep some distance to prevent its end effector from being damaged by the rising heat. When the stove is on high, the robot must stay an even greater distance. This simple scenario highlights a fundamental challenge for robot learning: contexts change, demanding different robot behaviors even when completing the same task.
@@ 1 Introduction
Traditional approaches to reward learning would treat these as separate tasks with distinct reward functions, even though they stem from the same underlying preference for safety. We argue this fails to capture a key observation: what often changes across contexts is not the human’s preference (e.g., valuing safety over efficiency), but the relative importance – or saliency – of reward features. In our example, the stove’s temperature doesn’t alter the human’s desire for safety, but rather modulates how much the robot’s stove distance matters in achieving that safety. Moreover, these contextual effects on feature saliency are independent of the specific task preference: whether a human generally prioritizes efficiency or safety, a cold stove renders the robot’s stove distance largely irrelevant. This aligns with hypotheses from cognitive science: humans are thought to construct value hierarchically, integrating features whose importance shifts with context o2021hierarchical . Even when high-level preferences remain stable, contextual changes can reshape feature saliency — a relationship we propose to explicitly model.
@@
Prior work in multi-task and meta IRL has focused on learning transferable representations across tasks yamada2022task ; yu2019meta , and may thereby capture some contextual effects. Yet these approaches can at best implicitly learn how context modulates features, often requiring substantial training data. Our key insight is that robots should instead explicitly learn how context affects feature saliency via targeted human feedback, rather than hoping these patterns will emerge with ever-growing amounts of data.
@@
We propose learning these relationships through calibrated features – representations that incorporate contextual effects on task reward features. Instead of learning separate reward functions for each context, our method learns how contextual elements (e.g. stove temperature) should affect task feature values. For example, when the stove is off, the learned calibrated feature should map any distance from the stove to the same value – reflecting that being 2 or 10 steps away makes no practical difference when there’s no heat to avoid. We learn calibrated features as a function of the base feature and the input state. Therefore, the representations capture how the inherent state context affects the feature saliency. By capturing these patterns of contextual influence individually for each feature, our method enables more efficient transfer to new scenarios, even if task preferences change.
@@
We introduce a novel type of human input to learn calibrated features: contextual feature queries – paired comparison queries that explicitly target how context affects feature saliency. A contextual feature query presents a user with two robot states, and asks them to select the context in which the robot better respects a particular feature. In Fig. 1, for the “laptop distance” feature, the user selects the first state with the empty coffee mug over the second state where the full coffee mug is dangerously close to the laptop. For the “stove distance” feature, the user labels the states as equivalent given in both states the stove is off, meaning stove distance is respected no matter the robot’s position. By focusing human feedback on these contextual relationships, we learn how different contexts reshape task features, enabling efficient adaptation to new tasks with minimal supervision.
@@
In summary, our contributions are three-fold. First, we introduce a formal framework for reward learning that explicitly models contextual effects through calibrated features, which can be reused in other tasks, enabling robots to efficiently adapt to new preferences. Second, we propose a method for learning context-calibrated features using targeted contextual feature queries. Finally, we empirically demonstrate that our approach improves sample efficiency and transferability across three complex reward learning tasks in robotic manipulation, outperforming multi-task baselines. A user study validates the ability of real people to respond to contextual feature queries, and uncovers the existence of diversity of contextual preferences, further emphasizing the importance of sample efficiency.
Figure 1: We propose learning representations that capture how context affects the saliency of reward features. (Step 1) The robot asks for contextual queries to learn context-calibrated features. (Step 2) Learned representations can be used to compose various contextually-adaptive downstream rewards.
2 Related Work
@@
Learning from Human Input. Recent advances in robot learning from human input primarily revolve around two paradigms: Inverse Reinforcement Learning (IRL) ng2000algorithms and Behavior Cloning (BC). BC directly learns a policy by mapping expert behavior to actions conditioned on the current state, making it susceptible to distribution shift. In contrast, IRL infers an underlying reward function that explains observed behavior and generalizes to new states. Various forms of human input for IRL have been explored, including demonstrations abbeel2004apprenticeship ; ziebart2008maximum , preferences christiano2017deep ; sadigh2017active , and corrections bajcsy2017learning ; jain2015learning . Beyond reward inference, prior work has also investigated representation learning from human input. Recent methods have aimed to learn transferrable reward representations, either implicitly learning encodings bobu2023sirl or explicitly learning reward features bobu2022inducing , though both without modeling contextual influences. Other recent work has explored using human input to learn safety constraints in the form of control barrier functions to specify unsafe robot behaviors (e.g., limiting a cup’s tilt angle) yang2024enhancing . Our work is complementary to this work. Instead of enforcing hard safety constraints, we provide more granular control by attenuating feature values rather than imposing strict boundaries.
@@
Modeling Context. Limited prior work explicitly models contextual effects on learned behaviors. Related work ghosal2023contextual models contextual effects as a binary switch determining which features are task-relevant. Rather than switching features on and off, we learn representations that account for the effect of context on saliency in a continuous, gradual way. Other work nishi2020fine ; sodhani2021multi structures multi-task rewards as a function of both features and context. These approaches assume explicit access to task context, in the form of metadata sodhani2021multi or discrete values nishi2020fine . In contrast, we don’t assume explicit access to the task context; instead, we learn which state elements compose the relevant context.
@@
Meta and Multi-Task Reward Learning. Multi-task and meta-IRL approaches aim to learn reward functions that transfer across tasks. Early multi-task IRL methods focused on clustering heterogeneous demonstrations and learn distinct rewards babes2011apprenticeship ; dimitrakakis2012bayesian ; choi2012nonparametric , which prevents information sharing across tasks. Following work has explored more complex reward representations with methods for learning shared task representations nishi2020fine ; yamada2022task . Developments in meta-IRL began with applications of MAML finn2017model to tabular settings xu2019learning and known task distributions gleave2018multi , followed by algorithms that introduce latent task variables for policy conditioning yu2019meta ; seyed2019smile . Recent approaches have addressed long-horizon planning chen2023multi and complex environments where maintaining a distribution over tasks is intractable wang2021meta . Prior work has shown that multi-task pre-training followed by fine-tuning can match or outperform meta-reinforcement learning while being more computationally efficient zhao2022effectiveness , therefore we compare our proposed method to a multi-task baseline yamada2022task . It is worth noting that unlike meta and multi-task reward learning methods, our approach does not require multiple training tasks.
3 Approach
3.1 Problem formulation
We are interested in teaching agents reward functions that adapt to context when executing a task. We denote the agent state as $s\in\mathcal{S}$, where $\mathcal{S}=\mathbb{R}^{d}$. We assume the agent has access to $M$ features $\phi(s):\mathbb{R}^{d}\rightarrow[0,1]$. [^1] We normalize base and calibrated features to $[0,1]$ for training stability, but in general they may lie in $\mathbb{R}^{+}$. These features represent useful task aspects (e.g. laptop distance), which can be hand-crafted ziebart2008maximum , learned bobu2022inducing , or even LLM-generated peng2024adaptive . A person’s desired task behavior naturally varies with contextual elements, like stove temperature or cup fullness. Rather than explicitly identifying which state elements constitute context, which is unrealistic in complex environments, we propose a more flexible approach that learns this relationship automatically. We directly model a person’s reward function as $R_{\theta}(\vec{\phi}(s),s):[0,1]^{M}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, where $\theta$ describes the function approximator parameters. The role of the learned reward $R_{\theta}$ becomes two-fold; it must capture which state elements comprise the relevant context and how these contextual elements should influence the reward. This enables adaptation to new contexts without requiring them to be explicitly labeled.
3.2 Structuring reward with calibrated features
@@
We propose modeling the effects of context using an intermediate representation on top of each feature. We present calibrated features as mappings from a base feature and state to a calibrated feature value, $\phi_{\psi_{i}}^{\prime}(\phi_{i}(s),s):[0,1]\times\mathbb{R}^{d}\rightarrow[0,1]$, where the $i$-th calibrated feature, a function approximator parameterized by $\psi_{i}$, corresponds with the $i$-th base feature. For notational convenience, we use $\phi_{\psi_{i}}^{\prime}(s)$ to denote $\phi_{\psi_{i}}^{\prime}(\phi_{i}(s),s)$, and $\vec{\phi}_{\psi}^{\prime}(s)$ to represent the vector of calibrated features computed separately for each base feature. The role of the $i$-th calibrated feature function is to capture which elements of the state should affect the saliency of the $i$-th base feature, and how. In our example (Fig. 1), the state may consist of robot joint angles, poses of objects (e.g., a cup, laptop, stove, etc.), and properties of those objects (e.g., stove heat or cup fullness). A learned calibrated feature would take as input the state and a base feature, such as the cup angle, and infer which state elements should condition the cup angle feature value (cup fullness) and how (saliency of cup angle is positively correlated with cup fullness). Importantly, we do not learn a function that outputs feature saliency, but one that conditions an input feature on the state context, producing a new, reshaped feature.
@@
We then learn reward on top of calibrated features, $R_{\theta}(\vec{\phi^{\prime}}_{\psi}(s)):[0,1]^{M}\rightarrow\mathbb{R}$, rather than as a function of base features and state. When the reward is linear, parameters $\theta$ maintain their traditional role quantifying the trade-off between features, while calibrated features capture contextual variation within a feature. Shown in Fig. 1, reward parameters may quantify the trade-off between staying away from hot objects and being careful with liquids near technology, while calibrated features reshape how each feature influences the corresponding higher-level concept, based on the current context.
3.3 Learning calibrated features

### v3

# Learning Contextually-Adaptive Rewards via Calibrated Features

@@ Abstract.
A key challenge in reward learning from human input is that desired agent behavior often changes based on context. For example, a robot must adapt to avoid a stove once it becomes hot. We observe that while high-level preferences (e.g., prioritizing safety over efficiency) often remain constant, context alters the saliency–or importance–of reward features. For instance, stove heat changes the relevance of the robot’s proximity, not the underlying preference for safety. Moreover, these contextual effects recur across tasks, motivating the need for transferable representations to encode them. Existing multi-task and meta-learning methods simultaneously learn representations and task preferences, at best implicitly capturing contextual effects and requiring substantial data to separate them from task-specific preferences. Instead, we propose explicitly modeling and learning context-dependent feature saliency separately from context-invariant preferences. We introduce calibrated features–modular representations that capture contextual effects on feature saliency–and present specialized paired comparison queries that isolate saliency from preference for efficient learning. Simulated experiments show our method improves sample efficiency, requiring 10x fewer preference queries than baselines to achieve equivalent reward accuracy, with up to 15% better performance in low-data regimes (5–10 queries). An in-person user study (N=12) demonstrates that participants can effectively teach their personal contextual preferences with our method, enabling adaptable and personalized reward learning.
robot reward learning, human preferences, representation learning
copyright: ccdoi: 10.1145/3757279.3788669journalyear: 2026isbn: 979-8-4007-2128-1/2026/03conference: Proceedings of the 21st ACM/IEEE International Conference on Human-Robot Interaction; March 16–19, 2026; Edinburgh, Scotland, UKbooktitle: Proceedings of the 21st ACM/IEEE International Conference on Human-Robot Interaction (HRI ’26), March 16–19, 2026, Edinburgh, Scotland, UKccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Learning from demonstrationsccs: Computing methodologies Cognitive robotics
Conceptual depiction of framework. Users first respond to contextual feature queries to teach calibrated features then train rewards on top of these calibrated features. Figure illustrates how the same three trajectories might be valued differently when contexts or user preferences change.
Figure 1. We learn representations that capture how context affects reward feature saliency. (Step 1) Contextual queries reveal when context changes feature importance: stove distance is better respected when the stove is off (top), whereas laptop distance is equally respected whether the cup is empty near the laptop or full but far from it (bottom). (Step 2) Calibrated features enable learning contextually adaptive rewards (high rewards in green and low rewards in red).
Conceptual depiction of framework. Users first respond to contextual feature queries to teach calibrated features then train rewards on top of these calibrated features. Figure illustrates how the same three trajectories might be valued differently when contexts or user preferences change.
1. Introduction
@@
Imagine a robot helping you in the kitchen. When the stove is off, the robot can pass freely above it to reach the counter. As the stove heats up, the robot should keep its end effector at a safe distance from the heat, and when the stove is on high, that distance should be even greater. This simple example highlights a core challenge for robot learning: contexts change, demanding different robot behaviors. Traditional reward learning frameworks would treat these scenarios as separate tasks with distinct rewards, even though they stem from the same underlying preference for safety. We argue that what actually changes across contexts is not the preference itself (e.g., safety over efficiency), but the relative importance–or saliency–of reward features. In our example, the stove’s temperature doesn’t alter the human’s desire for safety, but rather modulates how much the robot’s stove distance matters in achieving it. Critically, these contextual effects on feature saliency are often consistent across rewards and independent of specific preferences: whether a user prioritizes efficiency or safety, a cold stove renders stove distance irrelevant. This view aligns with cognitive science theories of hierarchical value construction (O’Doherty et al., 2020), and motivates the need for reusable representations that capture contextual effects on feature saliency, independent of task-specific rewards. Such representations would enable efficient learning of rewards that adapt across contexts.
@@
Predefined feature representations rarely capture contextual effects and are difficult to personalize, motivating approaches that learn representations from data. Multi-task and meta inverse reinforcement learning (IRL) methods aim to learn transferable structure by training shared representations across multiple tasks (Yamada et al., 2022; Yu et al., 2019). In principle, they could capture shared contextual effects–for example, learning that stove temperature consistently influences distance preferences across users’ reward functions. In practice, however, they must learn two things simultaneously: how context modulates feature saliency and how users weight different features. As a result, contextual effects are captured only implicitly, entangled with task-specific trade-offs, and disentangling the two requires large amounts of training data. Our key insight is that robots should instead explicitly model and learn how context affects feature saliency via targeted human feedback, rather than hoping these patterns will emerge with ever-growing data.
@@
We propose learning these relationships through calibrated features–representations that incorporate contextual effects on task reward features. Calibrated features explicitly capture how contextual elements (e.g., stove temperature) should affect task feature values. For example, when the stove is off, the learned calibrated feature should map any distance from the stove to the same value–reflecting that being 2 or 10 steps away makes no practical difference when there’s no heat to avoid. We learn calibrated features as a function of the base feature and the input state. Therefore, the representations capture how the inherent state context affects feature saliency. By capturing contextual effects individually for each feature, we learn contextually-aware representations that can be modularly reused and composed in new rewards.
@@
We introduce a novel type of human input for learning calibrated features, contextual feature queries. These paired comparison queries target how context affects feature saliency. Each query presents a user with two robot states, and asks them to select the context in which the robot better respects a particular feature. In Fig. 1, for the “stove distance” feature, the user selects the second state with the stove off over the first state where the robot is dangerously close to the hot stove. For the “laptop distance” feature, the user labels the states equivalent since in one state the cup is empty and the laptop safe regardless of the robot’s position, and in the other state the cup contains liquid but is held far from the laptop. By focusing human feedback directly on these contextual relationships, we can efficiently learn how context reshapes task features.
@@
In summary, our contributions are three-fold. First, we introduce a framework for learning representations that explicitly model contextual effects through calibrated features. Second, we propose a method for learning calibrated features using targeted contextual feature queries. Finally, we empirically show our approach improves sample efficiency in three complex reward learning tasks in robotic manipulation, outperforming multi-task baselines. A user study validates that real people can respond to contextual feature queries, and reveals diverse user contextual preferences, emphasizing the importance of sample efficiency for learning from individual input.
2. Background
@@
Learning from Human Input. Recent advances in robot learning from human input primarily revolve around two paradigms: Inverse Reinforcement Learning (IRL) (Ng and Russell, 2000) and Behavior Cloning (BC). BC learns a state-conditioned policy from expert actions, making it vulnerable to distribution shift, whereas IRL infers a reward function that explains observed behavior and generalizes to new states. Various forms of human input for IRL have been explored, including demonstrations (Abbeel and Ng, 2004; Ziebart et al., 2008), preferences (Christiano et al., 2017; Sadigh et al., 2017), and corrections (Bajcsy et al., 2017; Jain et al., 2015). Beyond reward inference, prior work has also investigated representation learning from human input. Existing methods aim to learn transferrable reward representations, either implicitly learning encodings (Bobu et al., 2023) or explicitly learning features (Bobu et al., 2022), though both without modeling contextual effects. Other recent work uses human input to learn safety constraints as control barrier functions to specify unsafe robot behaviors (e.g., cup tilt angle limit) (Yang et al., 2024). Our work is complementary; instead of enforcing hard safety constraints, we provide granular control by attenuating feature values.
@@
Modeling Context. Limited prior work explicitly models contextual effects on learned behaviors. Related work (Ghosal et al., 2023) models contextual effects as a binary switch determining which features are task-relevant. Rather than switching features on and off, we learn representations that capture the effect of context on saliency in a continuous, gradual way. Other work (Nishi and Shimosaka, 2020; Sodhani et al., 2021) structures multi-task rewards as a function of both features and context, assuming explicit access to task context either as metadata (Sodhani et al., 2021) or discrete values (Nishi and Shimosaka, 2020). In contrast, we don’t assume access to the context, instead learning which state elements compose the relevant context.
@@
Meta and Multi-Task Reward Learning. Multi-task and meta-IRL approaches aim to learn representations that transfer across tasks. Early multi-task IRL methods focused on clustering heterogeneous demonstrations and learning distinct rewards (Babes et al., 2011; Dimitrakakis and Rothkopf, 2012; Choi and Kim, 2012), limiting information sharing across tasks. Following work explored more complex reward representations with methods for learning shared task representations (Nishi and Shimosaka, 2020; Yamada et al., 2022). Developments in meta-IRL began with applications of MAML (Finn et al., 2017) to tabular settings (Xu et al., 2019) and known task distributions (Gleave and Habryka, 2018), followed by the introduction of latent task variables for policy conditioning (Yu et al., 2019; Ghasemipour et al., 2019). Recent work has addressed long-horizon planning (Chen et al., 2023) and complex environments where maintaining a distribution over tasks is intractable (Wang et al., 2021). Given prior work showing multi-task pre-training can match or outperform meta-RL while being more computationally efficient (Zhao et al., 2022), we compare our method to a multi-task baseline (Yamada et al., 2022). Finally, multi-objective RL (MORL) considers a setting like ours, modeling rewards as functions of base objectives, but typically focuses on Pareto-optimal policies (Moffaert and Nowé, 2014). State-of-the-art methods for learning MORL objectives (Mu et al., 2024) assume known training rewards, which is more restrictive than our baseline. Unlike all of these approaches, our method doesn’t require multiple training tasks.
3. Approach
3.1. Problem formulation
We are interested in learning representations that adapt to context for better downstream reward learning. We model our problem as a Markov decision process $(\mathcal{S},\mathcal{A},\mathcal{T},R)$, with states $s\in\mathcal{S}\subseteq\mathbb{R}^{d}$, actions $a\in\mathcal{A}$, transition $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$, and rewards $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. To express the reward function, we assume the agent has $K$ base features $\phi(s):\mathbb{R}^{d}\rightarrow[0,1]$ [^1] We normalize base and calibrated features to $[0,1]$ for training stability, but in general they may lie in $\mathbb{R}^{+}$., which represent useful task aspects (e.g., laptop distance) and can be hand-crafted (Ziebart et al., 2008), learned (Bobu et al., 2022), or even LLM-generated (Peng et al., 2024). We denote the feature set as $\Phi(s)=\{\phi_{i}(s)\}_{i=1}^{K}$, where $\phi_{i}(s)$ is the $i$-th feature.
We assume context is already captured in the state as elements $s_{c}\in s$ that affect the user’s desired behavior; for instance, stove temperature affects desired stove distance and cup fullness affects desired cup orientation. Since in general it’s unrealistic to assume the robot knows which state elements constitute context or how they affect behavior, we learn representations that infer these relationships directly from the state $s$. We formulate this problem as learning a representation $f_{\psi}(\Phi(s),s):[0,1]^{K}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{h}$, parameterized by $\psi$. For brevity, we denote this mapping as $f_{\psi}(s)$. The role of $f_{\psi}$ is two-fold; it must capture which state elements comprise the relevant context $s_{c}$ and how these contextual elements should influence the reward. We can then learn a reward function on top of the learned representation, $R_{\theta}(f_{\psi}(s)):\mathbb{R}^{h}\rightarrow\mathbb{R}$, parameterized by $\theta$, producing a reward that can adapt across contexts.
One way to learn these representations is through multi-task or meta-learning: train $N$ different reward functions $\{R_{\theta_{i}}\}_{i=1}^{N}$ that share a similar underlying representation $f_{\psi}(s)$. In this setup, the model must jointly learn task-specific parameters $\{\theta_{i}\}_{i=1}^{N}$ and shared representation parameters $\psi$. Disentangling shared contextual structure from task-specific preferences requires many diverse training tasks, making these approaches data-hungry and inefficient. Moreover, these methods learn representations that can’t be selectively reused or combined with new features in new rewards without retraining. We identify that end-user input can be more effective when directed toward explicitly describing contextual effects, rather than providing only general preference signals. Furthermore, if we learn feature-specific representations, these can be modularly reused and composed in downstream rewards.
3.2. Structuring reward with calibrated features
@@
We propose modeling the effects of context using an intermediate representation on top of each contextually affected feature. We present calibrated features as mappings from a base feature and state to a calibrated feature value, $\phi_{\psi_{i}}^{c}(\phi_{i}(s),s):[0,1]\times\mathbb{R}^{d}\rightarrow[0,1]$, where the $i$-th calibrated feature, a function approximator parameterized by $\psi_{i}$, corresponds with the $i$-th base feature. We denote $\phi_{\psi_{i}}^{c}(\phi_{i}(s),s)$ as $\phi_{\psi_{i}}^{c}(s)$ for notational convenience. The $i$-th calibrated feature serves to capture which elements of the state should affect the saliency of the $i$-th base feature, and how. A learned calibrated feature would take as input the state and a base feature, such as the cup angle, and infer which state elements should condition the cup angle feature value (cup fullness) and how (saliency of cup angle is positively correlated with cup fullness). Importantly, we do not learn a function that outputs feature saliency, but one that conditions an input feature on the state context, producing a new, reshaped feature. Because each calibrated feature depends only on its base feature and the state, it can be reused compositionally with new features as long as the state input remains the same.
Since our approach operates on semantically meaningful features, we assume that users can indicate which features should be contextually affected. This assumption is further discussed in Sec. 6. For a given task, we compose the representation $\Phi^{c}(s)=\{\phi_{i}^{c}(s)\}_{i=1}^{K}$ as a mix of calibrated and uncalibrated features, swapping in calibrated features when contextual effects exist,
(1) $\phi_{i}^{c}(s)=\begin{cases}\phi_{\psi_{i}}^{c}(\phi_{i}(s),s),&\text{if $c_{i}$},\\[6.0pt] \phi_{i}(s),&\text{otherwise},\end{cases}$
@@
where $c_{i}\in\{0,1\}$ is a binary indicator provided by the end user denoting if the $i$-th feature is contextually affected. We learn rewards on top of this representation, $R_{\theta}(\Phi^{c}(s)):[0,1]^{K}\rightarrow\mathbb{R}$. When the reward is linear, parameters $\theta$ maintain their traditional role quantifying the trade-off between features, while calibrated features capture contextual variation within a feature. Shown in Fig. 1, reward parameters describe the trade-off between staying away from hot objects and being careful with liquids near technology, while calibrated features reshape how each feature influences the corresponding higher-level concept, given the context.
3.3. Learning calibrated features